OpenStudy (anonymous):

WILL MEDAL AND FAN.. The numbers of turtles (t) in a lake after (n) months can be defined by the recursive sequence Tn=1.04Tn-1+3 and T0=100. How many turtles are in the pond after 3 months? Round to nearest whole number

OpenStudy (anonymous):

are you sure the formula is \[Tn=1.04Tn-1+3\] not \[Tn=1.04T0-1+3\]

OpenStudy (anonymous):

if the formula lime these it could rewrite: \[Tn=1.04Tn-1+3\] \[Tn=1.04Tn+2\]

OpenStudy (freckles):

\[T_n=1.04T_{n-1}+3?\]

OpenStudy (danjs):

T(n) = T(n-1) + 3 T(0) = 100 T(1) = 1.04T(0) +3 = 1.04*100 + 3 = 107 T(2) = 1.04T(1) + 3 = 1.04*107 + 3 =

OpenStudy (danjs):

and so on....

OpenStudy (danjs):

Recursion requires that you know the value of the term immediately before the term you are trying to find.

OpenStudy (danjs):

Find the T(2) from above, then T(3) = 1.04*T(2) + 3 = turtles in pond after 3 months.

OpenStudy (anonymous):

3.26 or 3 turtles in the pond? @danjs

OpenStudy (danjs):

No, T(2) means you use T(1) value T(3) means you use T(2) value... see the recursion? T(2) = 1.04*107 + 3 = 114.28 turtles T(3) = 1.04*114.28 + 3 = 121.85 T(3) rounded to nearest whole number = about 122 turtles

OpenStudy (danjs):

Remember after you get an answer to check if it is reasonable. The initial number of turtles, T(0) = 100 turtles. The function increases the amount each month by, 1.04 times the previous month and add 3. So after 3 months you would expect more than 100 turtles in the pond. Three turtles is not reasonable.